# A Survey of Deep Learning for Geometry Problem Solving

### v4

@@ Abstract
Geometry problem solving is a key area of mathematical reasoning, which is widely involved in many important fields such as education, mathematical ability assessment of artificial intelligence, and multimodal ability assessment. In recent years, the rapid development of deep learning technology, especially the rise of multimodal large language models, has triggered a widespread research boom. This paper provides a survey of the applications of deep learning in geometry problem solving, including (i) a comprehensive summary of the relevant tasks in geometry problem solving; (ii) a thorough review of related deep learning methods; (iii) a detailed analysis of evaluation metrics and methods; and (iv) a critical discussion of the current challenges and future directions that can be explored. Our goal is to provide a comprehensive and practical reference of deep learning for geometry problem solving to promote further developments in this field. We create a continuously updated list of papers on GitHub: https://github.com/majianz/dl4gps.
@@ 1 Introduction
As a core component of mathematical reasoning, Geometry Problem Solving (GPS) has been closely related to the field of education and the evaluation of mathematical ability of AI for many years Narboux et al. (2018). Due to the close connection between geometry problems and geometric diagrams, it has naturally become a typical representative of multimodal mathematical problems. Solving geometry problems in the format of educational exams not only requires artificial intelligence systems to understand and interpret geometric diagrams, but also requires strong logical reasoning and numerical calculation capabilities, which makes it an ideal scenario for evaluating the perception and reasoning capabilities of deep learning models. In recent years, the rise of Multimodal Large Language Models (MLLMs) has further promoted the development of this field, fully demonstrating the great potential of deep learning models in handling complex geometry-related solving and visual understanding tasks. The number of papers on deep learning for geometry problem solving grows approximately from 1 in 2018 to 110 in 2024 and continues to increase in 2025 (see Figure 5).
@@ 3.1.1 Encoder-Decoder Architecture
Knowledge Module. Some geometry problem solving systems integrate knowledge modules based on deep neural networks to more efficiently retrieve and apply knowledge and theorems in the field of geometry and verify the correctness of solutions. Knowledge Extractor and Integrator extracts and integrates geometric knowledge into the solution process. Some studies build geometric knowledge frameworks by constructing knowledge graphs. Fu et al. (2019) and Zhou et al. (2022) use BiLSTM to extract geometric relationships, and Tsai et al. (2021) uses Graph Convolutional Network (GCN) Kipf and Welling (2017) to embed knowledge graphs into vector space.  Xu et al. (2024b) and  Sharma et al. (2025) use CLIP and VISTA Zhou et al. (2024a) models to encode geometry problems to retrieve similar problems. Xiao et al. (2024a) builds a complete knowledge system through LSTM. Theorem Predictor is used to predict the geometric theorems required for the current solution step to derive a formal solution path. Guo and Jian (2022); Jian et al. (2023a) encodes the structural information of the formal language through GCN and uses the Sequence-to-Sequence (Seq2Seq) architecture Sutskever et al. (2014) of BiLSTM-GRU for theorem prediction. In addition, many studies use the Transformer-based Seq2Seq architecture for prediction Lu et al. (2021); Wu et al. (2024a); Zhang et al. (2024h), and some of these introduce the T5 model Yang et al. (2023); He et al. (2024b); Shengyuan and Xiuqin (2024). Zou et al. (2024) leverages DistilBERT Sanh (2019) to guide the training of theorem predictors. Answer Verifier is used to ensure the correctness of the solution logic. In addition to the traditional rule-based verification method Zhang et al. (2024f), Pan et al. (2025) introduces a pre-trained large language model Qwen et al. (2025) to verify the solution steps.
@@ 4.1.1 Performance-Based Metrics
Outcome-Based Metrics. Outcome-based metrics focus on measuring the correctness of the final inference results without considering the details of the reasoning process. Top-k Accuracy (Top-k Acc) and Pass@n (P@n) are two main metrics for measuring answer accuracy. Top-k Acc refers to the ratio of the first k predictions generated by the model that have at least one correct answer, while P@n refers to the ratio of the model that can correctly solve the problem at least once in a given n attempts. Other works also use outcome-based metrics such as F1 score Mishra et al. (2022b); Cheng et al. (2025b), maj@k Yue et al. (2024a), number of correct and wrong answers Dou et al. (2024), choice Zhang et al. (2023a) and competition scores. F1 score is a metric that takes into account both precision and recall; maj@k measures the ratio of cases that the model gets the correct answer through majority vote in k samples; the number of correct and wrong answers intuitively shows the task results; choice refers to the ratio of correctly selected answers from four candidate options of a multiple-choice question, with a random selection made if the model’s output does not match any of the given options; some studies also show model performance by showing SAT scores Seo et al. (2015) or IMO competition scores Trinh et al. (2024).
@@ 4.2.1 Manual Evaluation
In geometry problem solving, manual evaluation is relatively rarely used. It is done by human experts or annotators who directly check the output or reasoning process of the model. The core uses of manual evaluation include: first, evaluating the correctness of complex answers (e.g., judging whether $\frac{1}{\sqrt{2}}$ is equivalent to $\frac{\sqrt{2}}{2}$u et al. (2023); second, evaluating the interpretability of the reasoning process Sachan et al. (2017); Trinh et al. (2024). In addition, many studies will manually check the reasons for wrong and correct answers, which is also called a case study Lu et al. (2021); He et al. (2024a).
@@ Appendix A Geometry Problem Solving Datasets
In this section, we will further analyze various datasets for geometry problem solving. Table 1 and Table 2 provide a comprehensive summary of the datasets for geometry problem solving tasks from multiple perspectives, including dataset name, task type, geometry type, grade, problem source, whether it contains pictures, language, question type, rationale, training set and validation set size, test set size, and whether it is open source (if the data is open source, click the check mark to jump to the corresponding link).
There is a lack of geometry theorem proving data. Current academic research is mainly focused on geometric numerical calculation, while there is relatively little research on geometry theorem proving, and related data resources are also relatively scarce. There are many similarities between these two in terms of problem statement and mathematical knowledge Chen et al. (2022), but proof questions and calculation questions each have their own characteristics, and we should give them equal attention.
@@
There is a lack of solid geometry and analytic geometry data. The data currently used for geometry problem solving tasks is mainly concentrated in the field of plane geometry, while for other types of geometry problems, such as solid geometry Yu et al. (2021b) and analytic geometry Wu et al. (2023), there is still a lack of open-source training and testing data. A study points out that current solid geometry problems are often relatively simple and regular Xu et al. (2025b). Their diagrams usually only contain simple visual elements and rarely involve complex combinations of multiple geometric bodies, which to a certain extent limits the development of the field of solid geometry. Even in the field of plane geometry, high-quality data available for evaluation is relatively scarce recently.
There are limitations in data sources. Although these datasets are authentic and reliable, they are limited in quantity. Recently, due to the scarcity of real data and copyright protection issues, large-scale datasets are mostly constructed using data augmentation and program synthesis techniques Gao et al. (2025b); Pan et al. (2025). However, the data generated by these methods is insufficient in terms of authenticity, diversity, and quality, making it difficult to completely replace data from real scenarios.
There are deviations in language types and question types. In terms of language, existing data is mainly concentrated in English and Chinese, while real data for native speakers of other languages is extremely scarce Zhang et al. (2023b); Song et al. (2024), which not only limits the evaluation opportunities for relevant national exams, but also affects the fairness of the evaluation. In addition, the question types are mostly multiple-choice questions, which provide space for the model to guess the answer and are not conducive to the accurate evaluation of deep learning models.
@@
The datasets are also deficient in rationale. Most datasets lack detailed annotations of intermediate steps and reasoning processes Shi et al. (2024). Even if rationales exist, their formats are not uniform and granular enough to meet the goal of evaluating reasoning steps. In addition, the annotation format is mostly in natural language, which may not meet the needs of deep learning systems that reason in formal languages.
@@
Table 1: A summarization of geometry problem solving datasets for fundamental tasks and core tasks. Task: ER: Geometric Element Recognition, SR: Geometric Structure Recognition, DP: Geometric Diagram Parsing, DC: Geometric Diagram Captioning, SP: Semantic Parsing for geometry problem texts, RE: Geometric Relation Extraction, KP: Geometric Knowledge Prediction, TP: Geometry Theorem Proving, NC: Geometric Numerical Calculation. Type: P: plane geometry, S: solid geometry, A: analytic geometry. Question: MC: Multiple-Choice, NR: Numerical Response, FR: Free-Response, FB: Fill-in-the-Blank, YN: Yes-or-No, SA: Short-Answer, CQ: Classification Question. Rationale: nl: natural language. indicates that the dataset contains more than just geometry related content.
@@ Appendix B Other Geometry Tasks
In addition to geometry problem solving, there are other geometry-related tasks that have not been systematically summarized, which have similar fundamental tasks. More details for the datasets can be found in Table 3.
@@ B.1 Geometric Diagram Generation
This task is dedicated to generating high-quality geometric diagrams, aiming to facilitate a deeper understanding of geometry problems and related applications such as image editing, thereby providing strong support for the field of education.
Geometric Diagram Reconstruction. This task is one of the earlier works in the field of geometry. It aims to use existing simple sketches or preliminarily drawn images to reconstruct a clearer and more standardized complete image, thereby helping users to understand and visualize the image content more intuitively Yu et al. (2015). One of the key challenges is to reconstruct 3D geometry from 2D single line drawing images Xue et al. (2010, 2012); Yang et al. (2013), even if the input image is incomplete or inaccurate Zheng et al. (2015, 2016b, 2016a); Zou et al. (2016).
Geometric Text-to-Diagram. This task requires the system to be able to generate corresponding geometric diagrams from the natural language description of the geometry problem. This ability will greatly enhance the understanding of the solution system, enabling it to more accurately grasp the flexible and changeable geometric propositions Liu et al. (2012). In addition to traditional rule-based methods Janičić and Narboux (2021); Krueger et al. (2021); Trinh et al. (2024), some recent studies have begun to use deep learning technology to build related systems Zhengyu and Xiuqin (2023); Wang et al. (2025a); Cheng et al. (2025a). MagicGeoBench Wang et al. (2025a) collects a dataset containing 220 plane geometry problems in middle school mathematics exams, specifically for evaluating the performance of text-to-diagram geometry generation models.
In addition to the above methods, there are other techniques for generating geometric diagrams. Some tools support generating drawings through interactive ruler and compass construction, including GeoGebra [^1] https://www.geogebra.org and Geometer’s Sketchpad Scher (1999). Besides, the are some non-interactive methods for deriving such constructions Bertot et al. (2004); Itzhaky et al. (2013). In order to support more forms of geometric diagram generation, some studies have tried a wider range of methods to construct geometric diagrams, including techniques such as algebraic numerical optimization Gao and Lin (2004) and constrained numerical optimization Ye et al. (2020).
This task is also related to geometry problem solving. GeoX Xia et al. (2025) builds a pre-trained dataset with more than 120,000 plane geometry images, and tunes the visual encoder-decoder architecture through the mask auto-encoding scheme to obtain a visual encoder that fully understands geometric images. Some geometry problem solving work uses related methods to perform data enhancement on single-modal geometry problems and generate corresponding diagrams to obtain multimodal data Cai et al. (2024b); Zhao et al. (2024); Xiao et al. (2024b).
@@ B.2 Geometric Construction Problem
Geometric construction tasks, like geometry problem solving, are also part of educational exams. The goal is to use traditional ruler and compass construction methods to find an effective path to construct the target figure.
In recent years, some studies have tried to use deep learning systems to solve geometric construction tasks. In the online geometric construction game Euclidea [^2] https://www.euclidea.xyz,  Macke et al. (2021); Wong et al. (2023) uses Mask R-CNN He et al. (2017) to solve difficult geometric construction problems through a purely image-based method.  Mouselinos et al. (2024) converts the Euclidea problem into Python format and solves it with the help of a multi-agent framework based on large language models. This provides us with new ideas and inspires us to further explore the application potential of deep solution systems in cognitive fields such as planning and auxiliary line addition.
B.3 Geometric Diagram Retrieval
Before the widespread application of deep learning methods, the search problem of plane geometry diagrams has always been an important topic in the field of scientific research Liu et al. (2014a, b); Gan et al. (2016); Chen et al. (2016); Qu et al. (2016); Liu et al. (2016). With the development of computer technology, plane geometry retrieval may no longer be difficult in the era of deep learning. However, the retrieval of more complex solid geometry and irregular geometric figures may still be a direction worth studying.
@@ B.4 Geometric Autoformalization
Autoformalization is a subtask of theorem proving Li et al. (2024f). A small number of studies focus on automatically converting informal geometry problems and proofs into formal theorems and proofs that can be verified by machines. LeanEuclid Murphy et al. (2024) is a 173-problem geometric autoformalization dataset designed to test whether AI can understand mathematical problems and solutions written by humans and convert them into formal theorems and proofs.
@@
Table 4: A summarization of deep learning architecture for geometry problem solving system. Task: DP: Geometric Diagram Parsing, SP: Semantic Parsing for geometry problem text, ER: Geometric Element Recognition, DP: Geometric Diagram Parsing, SR: Geometric Structure Recognition, RE: Geometric Relation Extraction, TP: Geometry Theorem Proving, NC: Geometric Numerical Calculation. indicates the presence of the attention mechanism.

### v5

@@ Abstract
Geometry problem solving, a crucial aspect of mathematical reasoning, is vital across various domains, including education, the assessment of AI’s mathematical abilities, and multimodal capability evaluation. The recent surge in deep learning technologies, particularly the emergence of multimodal large language models, has significantly accelerated research in this area. This paper presents a survey of the applications of deep learning in geometry problem solving, including (i) a comprehensive summary of the relevant tasks in geometry problem solving; (ii) a thorough review of related deep learning methods; (iii) a detailed analysis of evaluation metrics and methods; and (iv) a critical discussion of the current challenges and future directions that can be explored. Our objective is to offer a comprehensive and practical reference of deep learning for geometry problem solving, thereby fostering further advancements in this field. We create a continuously updated list of papers on GitHub: https://github.com/majianz/dl4gps.
@@ 1 Introduction
As a core aspect of mathematical reasoning, Geometry Problem Solving (GPS) has long been closely tied to education and the assessment of mathematical proficiency in Artificial Intelligence (AI) systems (Narboux et al., 2018). Due to the close connection between geometry problems and geometric diagrams, it has naturally become a typical representative of multimodal mathematical problems. Solving geometry problems in the format of educational exams not only requires artificial intelligence systems to understand and interpret geometric diagrams, but also requires strong logical reasoning and numerical calculation capabilities, which makes it an ideal scenario for evaluating the perception and reasoning capabilities of deep learning models. In recent years, the rise of Multimodal Large Language Models (MLLMs) has further promoted the development of this field, fully demonstrating the great potential of deep learning models in handling complex geometry-related solving and visual understanding tasks. The number of papers on deep learning for geometry problem solving grows approximately from 1 in 2018 to 110 in 2024 and continues to increase in 2025 (see Figure 5).
@@ 3.1.1 Encoder-Decoder Architecture
Knowledge Module. Some geometry problem solving systems integrate knowledge modules based on deep neural networks to more efficiently retrieve and apply knowledge and theorems in the field of geometry and verify the correctness of solutions. Knowledge Extractor and Integrator extracts and integrates geometric knowledge into the solution process. Some studies build geometric knowledge frameworks by constructing knowledge graphs. Fu et al. (2019) and Zhou et al. (2022) use BiLSTM to extract geometric relationships, and Tsai et al. (2021) uses Graph Convolutional Network (GCN) Kipf and Welling (2017) to embed knowledge graphs into vector space.  Xu et al. (2024b) and  Sharma et al. (2025) use CLIP and VISTA Zhou et al. (2024a) models to encode geometry problems to retrieve similar problems. Xiao et al. (2024a) builds a complete knowledge system through LSTM. Theorem Predictor is used to predict the geometry theorems required for the current solution step to derive a formal solution path. Guo and Jian (2022); Jian et al. (2023a) encodes the structural information of the formal language through GCN and uses the Sequence-to-Sequence (Seq2Seq) architecture Sutskever et al. (2014) of BiLSTM-GRU for theorem prediction. In addition, many studies use the Transformer-based Seq2Seq architecture for prediction Lu et al. (2021); Wu et al. (2024a); Zhang et al. (2024h), and some of these introduce the T5 model Yang et al. (2023); He et al. (2024b); Shengyuan and Xiuqin (2024). Zou et al. (2024) leverages DistilBERT Sanh (2019) to guide the training of theorem predictors. Answer Verifier is used to ensure the correctness of the solution logic. In addition to the traditional rule-based verification method Zhang et al. (2024f), Pan et al. (2025) introduces a pre-trained LLM Qwen et al. (2025) to verify the solution steps.
@@ 4.1.1 Performance-Based Metrics
Outcome-Based Metrics. Outcome-based metrics focus on measuring the correctness of the final inference results without considering the details of the reasoning process. Top-k Accuracy (Top-k Acc) and Pass@n (P@n) are two main metrics for measuring answer accuracy. Top-k Acc refers to the ratio of the first k predictions generated by the model that have at least one correct answer, while P@n refers to the ratio of the model that can correctly solve the problem at least once in a given n attempts. Other works also use outcome-based metrics such as F1 score Mishra et al. (2022b); Cheng et al. (2025b), maj@k Yue et al. (2024a), choice Zhang et al. (2023a), number of correct and wrong answers Dou et al. (2024) and competition scores. F1 score is a metric that takes into account both precision and recall; maj@k measures the ratio of cases that the model gets the correct answer through majority vote in k samples; choice refers to the ratio of correctly selected answers from four candidate options of a multiple-choice question, with a random selection made if the model’s output does not match any of the given options; the number of correct and wrong answers intuitively shows the task results; some studies also show model performance by showing SAT scores Seo et al. (2015) or IMO competition scores Trinh et al. (2024).
@@ 4.2.1 Manual Evaluation
In geometry problem solving, manual evaluation is relatively rarely used. It is done by human experts or annotators who directly check the output or reasoning process of the model. The core uses of manual evaluation include: first, evaluating the correctness of complex answers (e.g., judging whether $\frac{1}{\sqrt{2}}$ is equivalent to $\frac{\sqrt{2}}{2}$Wu et al. (2023); second, evaluating the interpretability of the reasoning process Sachan et al. (2017); Trinh et al. (2024). In addition, many studies will manually check the reasons for wrong and correct answers, which is also called a case study Lu et al. (2021); He et al. (2024a).
@@ Appendix A Geometry Problem Solving Datasets
In this section, we further analyze various datasets for geometry problem solving. Table 1 and Table 2 provide a comprehensive summary of these datasets related to geometry problem solving tasks from multiple perspectives, including dataset name, task type, geometry type, grade level, problem source, presence of images, language, question format, rationale availability, sizes of training, validation, and test sets, as well as open-source status; a check mark indicates open-source datasets with links to the corresponding resources.
The current data for geometry theorem proving remains insufficient. Existing academic research predominantly centers on geometric numerical calculations, whereas studies on geometry theorem proving are relatively limited, and relevant data resources are still lacking. Despite sharing many similarities in problem formulation and underlying mathematical concepts (Chen et al., 2022), proof problems and calculation problems have distinct characteristics and challenges. Therefore, both types of geometry problems deserve equal attention.
@@
The current data for solid geometry and analytic geometry remains insufficient. Most datasets used in geometry problem solving tasks are concentrated in plane geometry, while data for other geometry types—such as solid geometry (Yu et al., 2021b) and analytic geometry (Wu et al., 2023)—remain limited. One study notes that existing solid geometry problems are often overly simple and regular (Xu et al., 2025b), with diagrams containing only basic visual elements and rarely involving complex geometric combinations, thereby restricting progress in this area. Even within plane geometry, high-quality evaluation datasets are still scarce.
The current data sources remain limited. While existing datasets are generally authentic and reliable, they are often small in scale. Recently, due to the shortage of real-world data and concerns over copyright, many large-scale datasets have been constructed via data augmentation or programmatic synthesis (Gao et al., 2025b; Pan et al., 2025). However, the synthetic data often falls short in terms of realism, diversity, and quality, making it difficult to serve as a full substitute for real data.
The current data coverage of language and question types remains limited. In terms of language, existing datasets primarily cover English and Chinese, while authentic data involving other native languages (Zhang et al., 2023b; Song et al., 2024) remains notably limited. This limits evaluation in the context of various national exams and reduces fairness. In terms of question types, most are multiple-choice, which allows models to guess answers and impairs accurate assessment of model reasoning ability.
@@
The current datasets remain lacking in rationale annotations. Most datasets do not provide detailed annotations of intermediate reasoning steps (Shi et al., 2024). Even when rationales are included, they often lack standardized formatting and sufficient granularity, falling short of the needs for evaluating step-by-step reasoning. Moreover, the rationale annotations are typically presented in natural language, which may not meet the needs of deep learning systems that operate in formal languages.
@@
Table 1: A summarization of geometry problem solving datasets for fundamental tasks and core tasks. Task: ER: geometric element recognition, SR: geometric structure recognition, DP: geometric diagram parsing, DC: geometric diagram captioning, SP: semantic parsing for geometry problem texts, RE: geometric relation extraction, KP: geometric knowledge prediction, TP: geometry theorem proving, NC: geometric numerical calculation. Type: P: plane geometry, S: solid geometry, A: analytic geometry. Question: MC: multiple-choice, NR: numerical response, FR: free-response, FB: fill-in-the-blank, YN: yes-or-no, SA: short-answer, CQ: classification question. Rationale: nl: natural language. indicates that the dataset contains more than just geometry-related content.
@@ Appendix B Other Geometry Tasks
In addition to geometry problem solving, some other geometry-related tasks, which have similar fundamental tasks, have not been systematically summarized. More details of the corresponding datasets can be found in Table 3.
@@ B.1 Geometric Diagram Generation
This task is dedicated to generating high-quality geometric diagrams. It aims to facilitate a deeper understanding of geometry problems and related applications such as image editing, thereby providing strong support for the field of education.
Geometric Diagram Reconstruction. This task is one of the earlier works in the field of geometry. It aims to use existing simple sketches or preliminarily drawn images to reconstruct a clearer and more standardized complete image, thereby helping users to understand and visualize the image content more intuitively (Yu et al., 2015). One of the key challenges is to reconstruct 3D geometry from 2D single line drawing images (Xue et al., 2010, 2012; Yang et al., 2013), even if the input image is incomplete or inaccurate (Zheng et al., 2015, 2016b, 2016a; Zou et al., 2016).
Geometric Text-to-Diagram. This task requires the system to be able to generate corresponding geometric diagrams from the natural language description of the geometry problem. This ability will significantly enhance the solution system’s understanding, enabling it to more accurately interpret geometric propositions presented in flexible and diverse forms (Liu et al., 2012). In addition to traditional rule-based methods (Janičić and Narboux, 2021; Krueger et al., 2021; Trinh et al., 2024), some recent studies have begun to use deep learning technology to build related systems (Zhengyu and Xiuqin, 2023; Wang et al., 2025a; Cheng et al., 2025a). MagicGeoBench (Wang et al., 2025a) provides a dataset of 220 plane geometry problems from middle school mathematics exams, designed to evaluate the performance of text-to-diagram geometry generation models.
In addition to the above approaches, various other techniques have been developed for generating geometric diagrams. Some tools, such as GeoGebra [^1] https://www.geogebra.org and Geometer’s Sketchpad (Scher, 1999), support interactive constructions using virtual ruler and compass operations to generate geometric diagrams. Additionally, non-interactive methods have also been proposed to automatically derive such constructions (Bertot et al., 2004; Itzhaky et al., 2013). To support more forms of geometric diagram generation, some studies have explored a wider range of methods to construct geometric diagrams. These methods include techniques like algebraic numerical optimization (Gao and Lin, 2004) and constrained numerical optimization (Ye et al., 2020).
This task is also related to geometry problem solving. GeoX (Xia et al., 2025) builds a pre-trained dataset containing more than 120,000 plane geometry images and tunes the visual encoder-decoder architecture using the mask auto-encoding scheme to obtain a visual encoder that fully understands geometric diagrams. Additionally, some geometry problem solving work uses related methods to perform data enhancement on unimodal geometry problems and generate corresponding diagrams to obtain multimodal data (Cai et al., 2024b; Zhao et al., 2024; Xiao et al., 2024b).
@@ B.2 Geometric Construction Problem
Geometric construction problems, similar to problems in geometry problem solving, are also part of educational exams. Such tasks aim to use traditional ruler and compass construction methods to find an effective way to construct the desired figure.
In recent years, some studies have tried to use deep learning systems to solve geometric construction problems. In the online geometric construction game Euclidea [^2] https://www.euclidea.xyz,  Macke et al. (2021); Wong et al. (2023) uses Mask R-CNN (He et al., 2017) to solve difficult geometric construction problems using a purely image-based method. Additionally, Mouselinos et al. (2024) converts the Euclidea problem into a Python format and solves it using a multi-agent framework based on LLMs. This provides us with new ideas and inspires us to further explore the application potential of deep learning systems in cognitive fields such as planning and auxiliary line addition.
B.3 Geometric Figure Retrieval
Before the widespread application of deep learning methods, the search problem for plane geometry figures had always been an important topic in the field of scientific research (Liu et al., 2014a, b; Gan et al., 2016; Chen et al., 2016; Qu et al., 2016; Liu et al., 2016). With the advancement of computer technology, plane geometry retrieval may no longer be challenging in the era of deep learning. However, retrieving more complex solid geometry and irregular geometric figures may still be a direction worth studying.
@@ B.4 Geometric Autoformalization
Autoformalization is a subtask of theorem proving (Li et al., 2024f). Autoformalization is a subtask of theorem proving (Li et al., 2024f). A few studies focus on automatically converting informal geometry problems and proofs into formal theorems and proofs verifiable by machines. LeanEuclid (Murphy et al., 2024) is a 173-problem geometric autoformalization dataset designed to test whether AI can understand mathematical problems and solutions written by humans and convert them into formal theorems and proofs.
@@
Table 4: A summarization of deep learning architectures for geometry problem solving system. Task: DP: geometric diagram parsing, SP: semantic parsing for geometry problem texts, ER: geometric element recognition, DP: geometric diagram parsing, SR: geometric structure recognition, RE: geometric relation extraction, TP: geometry theorem proving, NC: geometric numerical calculation. indicates the presence of the attention mechanism.